2. Consider a random variable with the following probability distribution: P(X=0) = 0.1, P(X=1) = 0.2, P(X=2) = 0.4, and P(X=3) = 0.3

a. Find P(X<=1)

b. Find P(1<X<=3)
Answer #1

(a) P(X <= 1) = P(X = 0) + P(X = 1) = 0.1 + 0.2 = 0.3

(b) P(1 < X <=3) = P(X = 2) + P(X = 3) = 0.4 + 0.3 = 0.7